OpenStudy (anonymous):

how to find the inverse function of f(x) = tan^-1 (10^x)

OpenStudy (dumbcow):

wait...the 10^x makes it a little different y = tan^-1(10^x) solve for x tan(y) = 10^x log(tan(y)) = x

OpenStudy (zzr0ck3r):

y = tan^(-1)(10^x) tan(y) = 10^x log((tany) = x

OpenStudy (zzr0ck3r):

notice we used log base 10, you could have used natural log but it would have been one more step..

OpenStudy (zzr0ck3r):

y = tan^(-1)(10^x) tan(y) = 10^x ln(tan(y)) = ln(10^x) ln(tan(y)) = x*ln(10) (ln(tan(y))/(ln(10)) = x

OpenStudy (zzr0ck3r):

so since we got to here log((tany) = x we "should" switch the x and y and then say y = f(x) so. f(x) = log(tan(x))

OpenStudy (anonymous):

thx for the detailed explanation :)